The table below lists the average distance R to the Sun and orbital period T of the first planets:
Distance Orbital Period
Mercury 0.39 AU 88 days
Venus 0.72 AU 225 days
Earth 1.00 AU 365 days
Mars 1.52 AU 687 days
(a) Calculate the average distance of Mercury, Venus and Mars to the Earth.
Which one of these planets is the closest to Earth on average?
(b) Calculate the average distance of Mercury, Venus and Earth to Mars.
Which one of these planets is the closest to Mars on average?
(c) What do you expect for the other planets?
1
Expert's answer
2020-06-22T11:20:02-0400

(a) The average distance of Mercury to the Earth:


"d_{ME}=d_E-d_M=1-0.39=0.61\text{ AU}."

The average distance of Venus to the Earth:


"d_{VE}=d_E-d_V=1-0.72=0.28\text{ AU}."


The average distance of Mars to the Earth:


"d_{MaE}=d_{Ma}-d_E=1.52-1=0.52\text{ AU}."

Venus is the closest planet to us.

(b) The average distance of Mercury to Mars:


"d_{MMa}=d_{Ma}-d_M=1.52-0.39=1.13\text{ AU}."

Venus to Mars:


"d_{VMa}=d_{Ma}-d_V=1.52-0.72=0.80\text{ AU}."

Mars to Earth: 0.52 AU.

(c) For other planets the distances will be much greater.